TInterval:

​(13.046,22.15)

x overbar =17.598

Sx= 16.01712719

n= 50

Express the confidence interval in the format that uses the​ "less than" symbol. Given that the original listed data use one decimal​ place, round the confidence interval limits accordingly.

____Mbps< μ <____Mbps

​(Round to two decimal places as​ needed.)
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (13.046 + 22.15) / 2 = 17.598

Margin of error = E = Upper confidence interval -   = 22.15 - 17.598 = 4.552

13.05 Mbps < < 22.15 MBPS